### v2

# Passivity-Centric Safe Reinforcement Learning for Contact-Rich Robotic Tasks

Abstract
Reinforcement learning (RL) has achieved remarkable success in various robotic tasks; however, its deployment in real-world scenarios, particularly in contact-rich environments, often overlooks critical safety and stability aspects. Policies without passivity guarantees can result in system instability, posing risks to robots, their environments, and human operators. In this work, we investigate the limitations of traditional RL policies when deployed in contact-rich tasks and explore the combination of energy-based passive control with safe RL in both training and deployment to answer these challenges. Firstly, we reveal the discovery that standard RL policy does not satisfy stability in contact-rich scenarios. Secondly, we introduce a passivity-aware RL policy training with energy-based constraints in our safe RL formulation. Lastly, a passivity filter is exerted on the policy output for passivity-ensured control during deployment. We conduct comparative studies on a contact-rich robotic maze exploration task, evaluating the effects of learning passivity-aware policies and the importance of passivity-ensured control. The experiments demonstrate that a passivity-agnostic RL policy easily violates energy constraints in deployment, even though it achieves high task completion in training. The results show that our proposed approach guarantees control stability through passivity filtering and improves the energy efficiency through passivity-aware training. A video of real-world experiments is available as supplementary material. We also release the checkpoint model and offline data for pre-training at Hugging Face.
I INTRODUCTION
In recent years, RL has earned increasing attention and success in addressing complex decision-making and control problems, especially in robotic applications [1]. From manipulation tasks to autonomous navigation, RL offers the potential to achieve unprecedented performance by learning optimal control policies. However, while RL excels in simulated environments, deploying these policies in real-world robotic systems remains a significant challenge due to the safety concerns [2]. This is particularly important in contact-rich tasks. Contact-rich tasks require the robots to interact with their environment safely, particularly in terms of contact force regulation, which is necessary to avoid damaging the environment, the robot, or the objects involved in the task. Although, there has been a significant effort to answer the contact-rich safety problem, the role of the control stability is often neglected within RL.
Stability is crucial for maintaining predictable and controllable robot behavior, especially when interacting with complex and unstructured surroundings [3] which can not be ignored in real-world robotic deployment [1]. For instance, an RL policy trained without considering system stability might generate control commands that cause high-frequency vibrations or unbounded energy accumulation. Such behaviors can damage the robot, compromise its performance, or pose safety risks to humans. Hence, ensuring that RL policies are inherently stable and safe is not only a theoretical concern but a practical necessity for robotics. Despite these risks, a substantial portion of RL research focuses primarily on maximizing task success rates or improving learning efficiency, with limited attention given to the stability of policy. This gap presents a significant barrier to real-world deployment, where stability and safety are not only desirable but essential [4].
Passivity-based control (PBC) is a fundamental method to prove the stability of physical interactions with the environment [5]. It is even more critical for kinematically redundant robots, as internal motions within the null space of the task Jacobian can result in an unintended increase in system energy, potentially leading to instability. Moreover, it is vital to monitor and restrict the amount of energy allowed to be stored and injected by the controller into the robot to ensure the passivity and safety of the system [6]. Energy tanks are widely used to achieve passivity-based control by restricting the amount of energy allowed to be stored in the system [7]. However, as mentioned above and pointed out in [8], the bounding of the energy alone is not sufficient to ensure the passivity of the system, but also the energy flow between the controller and robot has to be limited. We follow the recent works [9, 7, 10] with an augmented energy tank approach that also restricts the energy flow. Although energy tank-based passivity is studied extensively in the classical control literature, its inclusion in RL-based control learning is not sufficiently explored.
The inclusion of PBC in an RL system constitutes a significant gap in the RL literature. In this work, we aim to explore different modes of combination between RL and PBC, and study their implications. We term the standard RL approaches that do not include any passivity treatment as passivity-agnostic RL (Fig. 1.a). Then, we distinguish between the passivity-filtered approach, that imposes a passivity filter at the RL output; and the passivity-aware approach which aims to learn the passivity in the RL policy. In the former, the RL policy is trained traditionally, however, the filter analytically calculates the system energy and modifies the RL actions to stay passive in deployment. While it guarantees the passivity of the control, the RL agent remains unaware of the passivity. The latter introduces the energy awareness in the policy, however, it cannot guarantee the passivity as it is a black-box policy. Subsequently, we propose the passivity-ensured RL (Fig. 1.c) that includes both passivity guarantee and energy efficiency by combining the two approaches, as summarized in Table I.
Passivity-aware learning offers the possibility to obtain policies that solve tasks while providing benefits in terms of energy economy and long-term performance. Additionally, for systems with complex, nonlinear dynamics where analytical PBC design becomes intractable, learning provides a practical alternative that can discover passive policies without requiring complete system modeling. Work such as [11] and [12] use RL to enhance designed PBC systems. [11] uses RL to learn the parameters of the control system. [12] uses RL to dynamically adapt the size of an energy tank for changing tasks. To the best of our knowledge, only [13] employs the classical energy-tank based passivity approach during RL training by filtering the RL output and terminating the training episode when the energy tank is depleted. Although they train the system with an active filter, the RL agent is expected to learn the passivity only indirectly.
TABLE I: Method comparison
passivity- agnostic passivity- filtered passivity- aware passivity- ensured passivity guarantee energy efficiency
We advance the prior work by explicitly defining passivity-related constraints as part of the RL problem formulation. Furthermore, as stated in [8], solely limiting the total amount of energy is insufficient to ensure the passivity of the system. Hence, we augment the energy tank with a flow limit, unlike [13]. The main impact of this work lies in a systematic comparison of different approaches to combine PBC and RL. Furthermore, we formulate our passive-aware learning approach on a state-of-the-art safe RL framework to explicitly answer the contact-rich safety and passivity problems together. The safe RL concept has emerged to answer the safety concerns about the RL policies, which specify the safety requirements usually as constraints and aim to satisfy them while also solving the task [2]. We adopt a recent safe RL method that combines variable impedance control (VIC) and safety critic frameworks for contact rich tasks [14]. Leveraging this safe RL framework, we take a step towards a passivity-ensured safe RL policy for robotic contact-rich tasks that achieves dual safe and stable performance.
Lastly, we compare the four approaches of passivity-agnostic, passivity-filtered, passivity-aware and passivity-ensured RL extensively on a contact-rich interaction problem. The maze exploration task requires the robot to find the exit of a path by blindly touching the walls. Being contact-rich indicates a long exposure to physical contact. In fact, in this task, the contact is not only inevitable, but also fundamental to solve the task as the vision modality is not available. Our experimental study investigates the effect of passivity-based formalisms on the training/learning phase of the policy. Moreover, the performance in terms of task completion and energy expenditure of the different policies is compared to each other during deployment in simulation and the real world. The results confirm the advantages of the proposed approach in both guaranteeing the passivity and improving the energy usage while also safely solving a contact-rich task.
To summarize, the novel contributions of the manuscript are the following:
  1. 1.
    Extensive study of different RL-based passive control learning approaches (Table I), and their effect on both training and deployment phases, evaluated from safety, stability, energy efficiency and task completion perspectives.
  2. 2.
    Design and evaluation of the augmented energy tank-based passivity formulation in RL, with varying energy tank and energy flow constraints.
  3. 3.
    Introduction of passivity-ensured safe RL framework that simultaneously answers both control passivity and contact-rich safety problems.
[…]
Figure 1: System overview. The passivity-agnostic safe RL framework (a) constitutes the baseline method. Our contribution includes passivity-aware training (b) through energy constraints $C_{e}$, and passivity-ensured deployment (c) by adding a passivity layer to stabilize the RL action. We evaluate these three approaches through simulation and real-world experiments (d). Commonly, in all approaches, the RL agent observes the environment and outputs an action applied to the Cartesian VIC controller (Sec. II-B). The safe RL agent (e) uses a safety critic model to detect high risk and switch to the safer recovery policy. These models are pre-trained using offline data (f) consisting of random actions executed in multiple areas of the maze. The policy training (g) is completed by online training in simulation.
II Passive Safe Reinforcement Learning
The presented approach draws upon passive control and safe reinforcement learning fields to learn safe and stable policies for solving contact-rich tasks. As previously mentioned, it uses a passivity-based approach to observe and limit the energy stored in the system such that it can both: 1. impose constraints on the safe RL to train passivity-aware policies (Sec. II-D), and 2. achieve passivity-ensured control by a passivity layer running on the policy output (Sec. II-E). The former approach by itself does not guarantee passivity due to the incompleteness of data and the stochasticity of learning methods. The latter approach provides this guarantee, but can significantly reduce task performance by disrupting the task-oriented policy behavior. We combine both approaches to achieve good task performance while ensuring stability. A diagram of our work is shown in Fig. 1.
II-A Safe RL for contact-rich tasks
This section briefly introduces the contact-rich safe RL framework that forms the basis of our proposed method. The Safe RL network structure is shown in Fig. 1.e. Our RL agent consists of different components such as the safety critic model and recovery policy, in addition to the classic RL task policy. The safety critic is a separate network that predicts the potential risk of state-action pairs [15]. In the case of high risk, a recovery policy is activated to escape from risky areas [16]. The safety critic model and recovery policy are trained with the pre-collected offline data focusing only on the constraint violations. Then, all components including task policy are trained online with the maze exploration task. The policy output as action is given to the VIC, which controls the robot to achieve the desired actions. We use a variable impedance action-space for its advantages in physical interaction tasks [14, 17].
Safety can be formulated by introducing constraints to the RL policy in the form of a constrained Markov decision process (CMDP) [18]. A CMDP is expressed as a tuple $(\mathcal{S},\mathcal{A},R,P,\gamma,\mu,\mathcal{C})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $P$ is the state transition probability, $\gamma\in(0,1)$ is the reward discount factor, $\mu$ is the starting state distribution and $\mathcal{C}=\left\{\left(c_{i}:\mathcal{S}\rightarrow\{0,1\},\chi_{i}\in\mathbb{R}\right)\mid i\in\mathbb{Z}\right\}$ denotes the constraints that the agent must satisfy. For this specific formalism, we define the state space, action space, reward function, and safety constraints as below. All vectors are in the world reference frame unless specified.
  • State Space $\mathcal{S}$. The state space is designed as A 6-dimensional vector for safety critic network and recovery policy including the wrench $[f_{x},f_{y},f_{z},m_{x},m_{y},m_{z}]\in se^{*}(3)$ measured from an F/T sensor. Additionally, a 9-dimensional state vector is used, which extends this representation by incorporating the end-effector position [$p_{x},p_{y},p_{z}$]$\in\mathbb{R}^{3}$.
  • Action Space $\mathcal{A}$. Based on  [14], we design a 4-dimensional vector as action space, consisting of two end-effector displacement values $[\Delta p_{x},\Delta p_{y}]\in\mathbb{R}^{2}$ and 2 stiffness parameters $[k_{1},k_{2}]$. The action dimensions are bounded as listed in Table II. We keep the position $\Delta p_{z}$ and stiffness $k_{z}\in\mathbb{R}$ in the z-axis constant. The stiffness values are defined w.r.t. the motion vector $\Delta\mathbf{p}$ with $k_{1}\in\mathbb{R}$ along the motion vector, $k_{2}\in\mathbb{R}$ orthogonal to the motion vector and parallel to the ground plane, and $k_{z}$ normal to the ground plane.
  • Reward Function $R$. The agent is penalized for the Euclidean distance between the current position of the end-effector and the goal position multiplied by a constant $r_{pos}$. High-force collisions are penalized with $r_{col}$. A bonus is given for reaching the goal $r_{goal}$ to increase convergence speed. The reward multipliers used in our experiments are presented in Table II.
  • Constraints $\mathcal{C}$. The force constraint $c_{f}$ requires the measured force magnitude to stay below a threshold. We also propose two new constraints for passive control, described later in Sec. II-D.
In the following section, we describe how we apply the RL action with our low-level controller.
II-B Cartesian impedance control
The control scheme utilized in this work is a Cartesian VIC, for which the control torques of the impedance term $\boldsymbol{\tau}^{\top}_{\text{IC}}\in\mathbb{R}^{n}$ are defined in a quasi-static condition as
$\boldsymbol{\tau}^{\top}_{\text{IC}}=\boldsymbol{\tau}^{\top}_{\text{K}}-\boldsymbol{\tau}^{\top}_{\text{D}},\vspace{-0.1cm}$ (1)
where control torques $\boldsymbol{\tau}^{\top}_{\text{K}}$ are generated by the elastic wrench ${\mathbf{w}^{\text{0},\text{EE}}_{K}}^{\top}\in se^{\ast}(3)$, which can be expressed as
${\mathbf{w}^{\text{0},\text{EE}}_{K}}^{\top}=\begin{bmatrix}\mathbf{K}_{t}&\mathbf{K}_{c}\\ \mathbf{K}^{\top}_{c}&\mathbf{K}_{r}\\ \end{bmatrix}\Delta{\boldsymbol{\eta}}.\vspace{-0.05cm}$ (2)
Here, the infinitesimal body twist displacement is described by $\Delta{\boldsymbol{\eta}}\in se(3)$. The positive definite matrix $\mathbf{K}_{c}\in\mathbb{R}^{3\times 3}$ describes the decoupling between the stiffness values of the rotational ($\mathbf{K}_{r}\in\mathbb{R}^{3\times 3}$), and translational ($\mathbf{K}_{t}\in\mathbb{R}^{3\times 3}$) axes. In this work the values of translational stiffness matrix $\mathbf{K}_{t}$ are obtained by mapping the RL action stiffness matrix $\mathbf{K}_{a}=\operatorname{diag}(k_{1},k_{2},k_{z})$ (Sec. II-A) to the world frame as follows $\mathbf{K}_{t}=\mathbf{R}_{p}^{\top}\mathbf{K}_{a}\mathbf{R}_{p}$, given
$\mathbf{R}_{p}{=}\begin{bmatrix}\Delta p_{x}&-\Delta p_{y}&0\\ \Delta p_{y}&\Delta p_{x}&0\\ 0&0&1\\ \end{bmatrix}.$
The damping related torques are defined as $\boldsymbol{\tau}_{\text{D}}=\mathbf{D}\dot{\mathbf{x}}$, where the damping matrix $\mathbf{D}\in\mathbb{R}^{6\times 6}$ is formed as in [19] and $\dot{\mathbf{x}}\in se(3)$ is the Cartesian velocity of the end-effector. In the following subsection, we describe the passivity-based approach for guaranteeing the stability of our controller.
II-C Passivity-based approach for variable stiffness control
As mentioned in Section I, passivity is a necessity to facilitate physical interactions between a robot and its environment. For an impedance-controlled robot, the energy stored in the system can be expressed as the sum of the storage functions of the controller and robot $S=S_{\text{Ctrl}}+S_{\text{Rob}}\in\mathbb{R}$  [20]. It was shown in  [5] that since $S_{\text{Rob}}$ is physically bounded, it is also passive. Therefore, the system is passive if and only if the controller described by $S_{\text{Ctrl}}$ is passive. Hence, it is a necessity to observe and bound the energy injected by the controller at the power port $(\boldsymbol{\tau}_{\text{Ctrl}},\mathbf{\dot{q}})$, which is expressed by
$\dot{S}_{\text{Ctrl}}+P_{\text{IC}}=0.\vspace{-0.1cm}$ (3)
Here, $P_{\text{IC}}$ is expressed as the sum of the energy flows resulting from the spring $P_{\text{K}}=\mathbf{w}^{\text{0},\text{EE}}_{K}\dot{\mathbf{x}}\in\mathbb{R}$ and damping $P_{\text{D}}=\dot{\mathbf{x}}^{\top}\mathbf{D}\dot{\mathbf{x}}\in\mathbb{R}$. As mentioned previously in Sec. I, one can ensure the passivity of the overall system via the augmentation of the storage function $S_{\text{Ctrl}}$ by an energy tank $E_{\text{T}}\in\mathbb{R}$ that is bounded by the upper and lower limits $\overline{E}_{\text{T}}$/$\underline{E}_{\text{T}}\in\mathbb{R}$. Thus, the energy flow at the power port $(\boldsymbol{\tau}_{\text{Ctrl}},\mathbf{\dot{q}})$ can be rewritten as follows:
$\dot{S}_{\text{Ctrl}}+\dot{E_{\text{T}}}+P_{\text{D}}\leq 0,\vspace{-0.1cm}$ (4)
where the energy flow at the tank in this work is defined as $\dot{E}_{\text{T}}=-P_{K}$. The system passivity requires that the tank condition $\underline{E}_{\text{T}}\leq E_{\text{T}}{+}\dot{E_{\text{T}}}\leq\overline{E}_{\text{T}}$ to always hold.
However, as pointed out in [8], satisfying the tank condition is not enough to ensure the passivity of a system, but it is also a necessity to restrict the energy flow $\dot{E_{\text{T}}}$ between the controller and robot. This can be achieved by introducing a flow limit $\underline{\dot{E}}_{\text{T}}\in\mathbb{R}$, and ensuring that the flow condition $\underline{\dot{E}}_{\text{T}}\leq\dot{E}_{T}$ always holds.
In sections II-D and II-E, we go into more detail on how we incorporate energy tank and energy flow conditions into the learning and deployment phases, respectively.
II-D Passivity-aware safe RL
This section describes how we add passivity awareness to the RL policy. As described in Section II-A, the safe RL framework achieves safety through constraint definitions. We extend the CMDP formulation by including the passivity constraints so that the safety critic also learns the concept of passivity. As illustrated in Fig. 1.b, this information is provided as the energy constraints ${C_{e}}\subset\mathcal{C}$, including a budget constraint ${c_{eb}\in{C_{e}}}$ and a flow constraint ${c_{ef}\in{C_{e}}}$, which correspond to the energy tank and energy flow conditions.The budget constraint ${c_{eb}}$ is introduced to make the agent aware of the amount of energy it has available for adapting the Cartesian stiffness $\mathbf{K}$ during the task. The constraint is triggered whenever $E_{\text{T}}+\dot{E_{\text{T}}}\leq\underline{E}_{\text{T}}$, which indicates that the lower bound of the energy tank is reached.
Whereas the purpose of introducing the energy flow constraint ${c_{ef}}$ is to make the agent aware of how much energy it is allowed to take out of the energy tank at any given time for a specific task. Therefore ${c_{ef}}$ is triggered whenever the draining energy flow of the energy tank exceeds the defined threshold $\dot{E}_{\text{T}}<\underline{\dot{E}}_{\text{T}}$. RL training episodes are terminated when a constraint happens. This makes it easy to end an episode with both of the energy constraints, hindering the training process. Thus, we experiment with different combinations of these constraints as detailed in Sec. III. Defining the passivity information as constraints means that the training data includes this information. This data is used to train the safety critic, which learns the risk related to passivity. Therefore, we expect the passivity-aware RL agent to learn to use the energy more economically, in means of cumulative energy spent and instant energy flow. However, learning a regression model does not guarantee that the risk is always predicted correctly. Thus, we recommend deploying these models with the added passivity layer for passivity-ensured control, as described in the following section.
II-E Passivity-ensured control
We employ a passivity layer at the output of the RL agent to guarantee the control passivity, as illustrated in Fig. 1.c. The passivity layer contains two filtering actions to satisfy the energy budget and energy flow conditions defined in Section II-C. We ensure the flow condition [8] by limiting the rate at which the controller can inject energy into the system through the introduction of the scaling variable
$\alpha=\begin{cases}\frac{\underline{\dot{E}}_{\text{T}}}{\dot{E}_{T}}&\text{if }\dot{E}_{T}<\underline{\dot{E}}_{\text{T}}\leq 0\\ 1&\text{otherwise}.\vspace{-0.1cm}\end{cases}$ (5)
It is defined as the ratio between the maximal allowed energy flow $\underline{\dot{E}}_{\text{T}}$ and the originally calculated flow $\dot{E}_{T}$, resulting in a new formulation of the energy flow
$\dot{E_{\text{T}}}^{\prime}=\begin{cases}\alpha\dot{E_{\text{T}}}&\text{if }\underline{E}_{\text{T}}\leq E_{\text{T}}+\dot{E_{\text{T}}}\leq\overline{E}_{\text{T}}\\ 0&\text{otherwise}.\end{cases}\vspace{-0.1cm}$ (6)
Integrating the described energy tank dynamics for the VIC, the change of the Cartesian stiffness can be defined as
$\mathbf{K}(t)=\begin{cases}\mathbf{K}(t{-}1)&\text{if }E_{\text{T}}+\dot{E_{\text{T}}}^{\prime}\leq\underline{E}_{\text{T}}\\ \mathbf{K}(t)&\text{otherwise}.\\ \end{cases}\vspace{-0.1cm}$ (7)
In the event that the energy tank becomes depleted, $\mathbf{K}$ is hindered from increasing further, thereby keeping $\mathbf{K}$ constant. This results in a standard Cartesian impedance controller with constant gains. However, when $E_{\text{T}}>\overline{E}_{\text{T}}$ becomes true through a decrease of $\mathbf{K}$, the change of stiffness is applied but without storing more energy in the tank. In order to enforce the energy flow constraint, $\alpha$ is incorporated in the torques generated by the impedance control term:
${\boldsymbol{\tau}}^{\top}_{\text{IC}}=\alpha{\mathbf{J}}^{\top}(\mathbf{q})\big{(}{\mathbf{w}^{0,\text{EE}}_{K}}^{\top}+\mathbf{D}\dot{\mathbf{x}}\big{)}.\vspace{-0.1cm}$ (8)
These filtering actions are necessary to guarantee passive behavior. However, reconsidering Fig. 1.c, the RL agent perceives the low-level control and the passivity layer as part of the environment, because these are not part of the RL-based decision-making. Thus, any post-processing applied to the RL action decreases the transparency of the system, i.e., the outcome of the same action may be different depending on the filter variables. For this reason, we combine the passivity layer with passivity-aware policies and evaluate the effects of each component in our experiments.
III Evaluation
We evaluate our method through the contact-rich maze exploration task. In the following evaluation experiments, we aim to study: 1. how passivity constraints affect RL policy training and performance; 2. how to guarantee both safe and stable performance while executing the task efficiently. To do so, we conduct three types of experiments. First, we train the agent with different energy budget and energy flow constraints, comparing them to the passivity-agnostic baseline (Sec. III-B). Second, we deploy the learned models with the passivity layer (Sec. III-C), which guarantees control stability by limiting agent actions. Third, we validate our approach through real-world experiments (Sec. III-D).
Figure 2: Experiments setup in MuJoCo. Right top: the robot explores the turn area with contact. Right bottom: three sphere obstacles.
III-A Experimental design
The experimental task is a maze exploration task where the robot has to explore an unknown maze blindly. Thus, it relies on the force sensing for finding its way out, making it a contact-rich task, i.e., the robot does not avoid the contact and it has to regulate the exerted wrenches for safety. During this process, the controller dynamically adjusts Cartesian stiffness and pose in response to disturbances, such as maze walls and movable obstacles represented by three simulated sphere balls in the middle of the maze.
[…]
TABLE II: Experiment parameters
Reward multipliers $r_{pos}=-400$, $r_{col}=-250$, $r_{goal}=1000$ Action space $\Delta p_{x},\Delta p_{y}\in$ [-3, 3] cm, $k_{1},k_{2}\in$ [300, 1000] N/m Control parameters $k_{z}=$ 750, $\mathbf{K}_{r}=\operatorname{diag}(100,100,0)$, $\mathbf{K}_{c}=\mathbf{0}$ N/m Training hyperparameters learning rate: $3\times 10^{-4}$, $\gamma$: 0.9, $\gamma_{safe}$: 0.85, $\epsilon_{risk}$:0.65
Figure 3: Passivity-aware policy training experiment results. We present the total numbers of success and violation in the first two plots. The third plot shows the ratio of mean success to mean violation.
III-B Passivity-aware policy training
As mentioned previously, this section focuses on how the passivity constraints effect the training of a RL policy, by comparing them to each other as well as to the passivity-agnostic approach that is solely subjected to the contact-force constraint as shown in Fig. 1. More specifically the remainder will focus on energy budgets that are within the range of $4{-}10$ and flow limits are from $-0.7$ to $-0.3$, which were found based on preliminary trials. The results of the training experiments are summarized as means and standard deviations of 10 seeds in Fig. 3. More specifically, it shows the amount of success in terms of task completion, the type and number of constraint violations which let to the task termination as well as the ratio between the amount of successfully completed runs and violations.
Firstly, it can be seen that the tasks success is directly correlated with the amount of constraints introduced to the CMDP. Secondly, the introduction of $c_{eb}$ has a direct effect on the amount of force violation, which overall decreases as the energy budget gets smaller. Furthermore, the introduction of $c_{ef}$ results in that the force constraint are no longer violated. The only outlier is the Eb10, which outperforms the agnostic run, even though it violates both the budget and force constraint. Thirdly, it can be seen that the introduction of $c_{ef}$ has a more aggressive effect on the task success than the introduction of only $c_{eb}$. This is due to the fact that the constraint on the flow hinders the agent to make aggressive changes in its stiffnesses, resulting in an overly conservative approach which hinders the agent to explore the maze. Hence, generally it can be said that there is a consistent increase in constraint violations as the budget size decreases and harsher flow limits are applied.
Figure 4: These plots visualize the effect that the changes of the stiffnesses $k_{t_{x}},k_{t_{y}}\in\mathbf{K}_{\text{T}}$ during training, have on the energy flow $\dot{E_{\text{T}}}$ and consequently on the energy stored in the energy tank $E_{\text{T}}$.
Figure 5: Plots illustrate the effect that the introduction of the energy tank constraint has on the agents energy expenditure at different stages during training.
During training the agent regulates the stiffness of the impedance controller (Sec. II-B). This has a direct effect on how the energy flow $\dot{E_{\text{T}}}$ and the energy tank budget $E_{\text{T}}$, which is visualized in Fig. 4. More specifically, it is shown that whenever the agent increases the stiffness of springs during its motion, it results in $\dot{E_{\text{T}}}<0$ as long as $\underline{E}_{\text{T}}\leq E_{\text{T}}<\overline{E}_{\text{T}}$ holds as described in Sec. II-C.
Furthermore, it can be observed that the greater the increase in stiffness during a particular motion the greater effect it has on $\dot{E_{\text{T}}}$ and consequently on $E_{\text{T}}$. This behavior can be directly related to the definition of $\dot{E_{\text{T}}}$ (Sec. II-C). During training, one can observe how the agent becomes more aware of how the change of $\mathbf{K}$ effects either $\dot{E_{\text{T}}}$ and/or $E_{\text{T}}$ depending on the constraints it is subjected to. The impact that $C_{e}$ has on passivity-aware policy training can be seen in Fig. 5, showing the agent’s energy expenditure at two different stages during training. Where the red graph representing the energy expenditure and the energy flow at the port of the energy tank in the first episode, and the blue graph represents the agent’s energy-related behavior after 100 episodes. When comparing both graphs it is evident that the agent has learned across these 100 episodes to act more conservatively, resulting in reduced and more evenly distributed energy expenditure. Hence, it can be seen how introducing the energy constraints $C_{e}$ have a direct impact on the agent’s behavior during training.
TABLE III: Results of model tests. Performance evaluation over 100 episodes for each of three randomly selected seeds.
Run config. Violations Successes
Force Tank Flow $\mu\pm\sigma$
agnostic 0 0 0 $100\pm 0$
Eb4 0 26 0 $91.3\pm 3.9$
Eb8 2 0 0 $99.3\pm 0.9$
Eb8-Ef7 0 7 0 $62\pm 40.1$
Figure 6: Policy performance in deployment experiments. The bars show how the models (agnostic, Eb4, Eb8, Eb8-Ef7) perform in terms of successes and violation with and without the passivity layer in simulation for 100 episodes, subjected to the same constraints $\overline{E}_{\text{T}}{=}6$, $\dot{\underline{E}_{\text{T}}}{=}-0.5$.
After training, we selected the policies for agnostic, Eb4, Eb8, Eb8-Ef7 to test the model in representative experiments. All policies are evaluated over 100 episodes under the same constraint settings used during training. In the testing results, shown in Table III, similar to the training results shown in Figure 3, the agent achieves the highest success rate when it is agnostic to $C_{e}$ and can therefore operate more exploratory during its runs. Furthermore, looking at Eb8 and Eb4, a more conservative energy tank size results in more tank violations. However, compared to the training results, the Eb8-Ef7 run obtained less successes as it has harsher constraints during results. Here, Eb4 obtains high success, even though it had fewer successes than Eb8-Ef7 in training. From this it can be concluded that Eb4 learned a better task policy, although it was delayed by its limited energy budget.
[…]
Figure 7: (a) Energy budget progress during passivity-ensured deployment experiments. The curves indicate the mean of 300 runs (100 per seed) and the shades indicate the standard deviation. (b) Histogram of the run times, i.e., the number of runs that finished in time duration intervals.
[…]
Figure 8: Snapshots and result analysis in real-world experiments. (a) the overview of maze exploration task. (b) closeup when robot moves along with contact force. (c) failure case agnostic but non-passive behavior at the turnaround area. (d) shows the energy budget consumption among different setups, demonstrating our method outperforms in terms of passivity and energy efficiency.
III-C Passivity-ensured deployment
In the deployment experiments, we evaluate the effect of the passivity layer on the trained models, by comparing the performances of the passivity-agnostic baseline and various passivity-aware approaches in combination with and without the passivity layer. Where the passivity-agnostic approach with the passivity layer is called passivity-filtered and the passivity-aware approaches with the filter become passivity-ensured. In these experiments, we do not retrain the models but impose the passivity layer on the output of the RL agent (Fig. 1), where the layer is with uniform parameter settings: an initial energy budget of $\overline{E}_{\text{T}}{=}6$ and an energy flow limit of $\dot{\underline{E}_{\text{T}}}{=}-0.5$.
The results in Fig. 6 showcase the task successes and violations. Overall, the results labeled as “with passivity layer” outperform the baseline labeled as “no passivity layer”. These results reveal the risks of deploying models without passivity guarantees, which potentially leads to system instabilities and unsafe behaviors despite good performance in model tests shown in Table III.
Specifically: 1. The agnostic policy performs the worst when no passivity filter is applied, as it consistently violates the flow constraint. Although it achieves a 100% success rate when evaluated without considering energy constraints shown in Table III, which highlights the necessity of incorporating passivity during training. 2. The success rate improves significantly when the passivity filter is applied, underscoring the dual benefits of passive-aware policies and the passivity layer, ensuring safer and more stable deployments. 3. Violations dropped significantly with the passivity layer, with only Eb8 had two force violations that is almost not visible in Fig. 6. Note that no energy constraint violations occured, as the passivity layer ensures system passivity.
We also present the energy usage of each method in Fig. 7.a to further demonstrate the benefit of passivity-aware training. Even though all of agnostic, Eb8 and Eb4 show full success in the tests (Fig. 6), analysis of the energy budget progress in these runs reveals that the agnostic model depletes its budget more rapidly. This rapid depletion reduces its likelihood of success in more complex task instances. On the other hand, Eb4 demonstrates the best success and energy efficiency, despite its delayed learning shown in Fig. 3. Furthermore, the histogram in Fig. 7.b shows number of runs distribution on completion time. It is clear that both Eb8 and agnostic agents finish the task around 30s in majority of runs, while some runs take longer due to getting stuck in the maze. This highlights that conservative agents (Eb4, Eb8-Eb7) take slightly longer time to finish the maze on average.
III-D Real-world experiments
The trained models are deplyed in the real world to complete our validation on a Franka Panda robot arm with a flange mounted at the end effector with an ATI45 FT sensor to percept external force and torque as shown in Fig. 8 upper, with our passivity-ensured impedance controller (frequency 1000Hz). We run the best models of each setup 3 times in the real world without fine-tuning. The robot can finish the task without violating force constraints and energy constraints in each experiment, which validates our method’s effectiveness. However, failure case happened in the passivity-agnostic model. Specifically, the Fig. 8 lower shows all models’ energy expenditure over time steps. First, we can observe that the agnostic consumes energy most aggressively although keep passive due to passivity layer, while passivity-aware models maintain lower energy consumption rates. Notably, Eb4 achieves the most efficient energy usage among all models. This suggests that training with tighter energy constraints leads to more energy-efficient behaviors that transfer well to real-world deployment. The overall trend validates that our passivity-aware approach successfully learns energy-efficient policies that retain effectiveness in real-world settings. The video is available as supplementary material.
III-E Discussion
During training, the agent becomes more energy-conservative under passivity constraints $C_{e}$. These constraints also reduce task success rates by limiting exploration, especially early in learning, suggesting constrained cases need more training than unconstrained ones. However, overly restrictive constraints can prevent learning entirely if the task becomes physically infeasible.
Deployment results in Table III show that passivity cannot be guaranteed without a passivity layer, regardless of whether the agent was trained with passivity awareness. While the passivity layer improves task completion and reduces violations for all agents, those trained with energy constraints ($C_{e}$) demonstrate better energy efficiency.
The correlation between the energy usage trends in simulation (Fig. 7.a) and real-world experiments (Fig. 8 .d) demonstrates the successful transfer of learned energy-efficient behaviors. In both cases, the agnostic model shows the steepest decline in energy budget, while Eb4 maintains the most conservative energy usage. This consistent pattern validates that the energy-aware policies learned in simulation maintain their characteristics when deployed on physical hardware. The real-world results also confirm that tighter energy constraints during training (Eb4) lead to more efficient policies that better manage the energy budget in both simulated and real environments. However, the absolute rate of energy consumption differs between simulation and reality due to modeling discrepancies and real-world friction effects.
IV Conclusion and future work
In this letter, we investigated the safety and stability of the robotic system under multiple constraints for a contact-rich task, demonstrating that the traditional RL approach can not guarantee passivity. Furthermore, we combined an augmented energy tank-based passivity formalism in a safe RL framework. We employed the passivity constraints in both training and deployment of the RL model. The proposed method completed the task without any force violation while ensuring system passivity.Additionally, passivity-aware policies showed improved energy efficiency in deployment, even though they converged later in the training. A set of experiments both in simulation and in the real world on a 7-DoF Franka robot arm validated the effectiveness of the proposed method. We are planning to explore self-adaptive energy tank-size and flow constraints in the future works.